OpenStudy (anonymous):

Please HELP. Test tomorrow! Convergence of Series I have sum(2 to inf) of 1/(ln(n))^2 I'm fairly certain it diverges. I'm not sure what test to use to show that it diverges though. Any thoughts?

OpenStudy (tkhunny):

Ponder the relationship between \[f(x) = ln(x)\] and \[g(x) = \sqrt{x}\]

OpenStudy (anonymous):

I think just compare it to 1/n....

OpenStudy (anonymous):

what were you saying @tkhunny ?

OpenStudy (anonymous):

I don't get it... sqrt(x)?

OpenStudy (tkhunny):

\[ln(x) < \sqrt{x}\] For x > 1, they are both positive, so... \[\left(ln(x)\right)^{2} < x\]. Are we getting anywhere?

OpenStudy (tkhunny):

Whoops, for x > 2 and x an integer, they are both greater than 1. That's better.

OpenStudy (anonymous):

guess we're saying the same thing... compare 1/(lnx)^2 to 1/x...

OpenStudy (tkhunny):

Well, we may both be saying it, but neither of us has has quite proved it. What say you, cosmonautics, is the inuendo good enough or do we need a better demonstration?

OpenStudy (anonymous):

well, you showed it... if (lnx)^2 is less than x then 1/(lnx)^2 is greater than 1/x 1/x diverges...